After going to the store, they had a special on paint with 7% color included. However, you don't want that bright of a color, but rather 3% color. How much 7% paint do you buy to mix with plain white paint to get your 3% mixture? The total amount of piant needs to equal 5.94 liters.
I need the amount of 7% paint and the amount of white paint.

Answer by ankor@dixie-net.com(22740)

After going to the store, they had a special on paint with 7% color included. However, you don't want that bright of a color, but rather 3% color. How much 7% paint do you buy to mix with plain white paint to get your 3% mixture? The total amount of paint needs to equal 5.94 liters.
I need the amount of 7% paint and the amount of white paint.
:
Let x = amt of 7% paint required
:
It tells us the total has to be 5.94 liters, therefore:
(5.94 - x) = amt of white paint required
:
Write a % equation in decimal form:
.07x = .03(5.94)
:
.07x = .1782
:
x = .1782/.07
:
x = 2.5457 liters of 7% paint required
:
5.94 - 2.5457 = 3.3943 liters of white paint
:
:
Check solution by finding the amt of color in 7% paint and the mixed paint
.07 * 2.5457 = .1782
.03 * 5.94 = .1782